# October   ||   2024 AI-Driven Control of Bioelectric signalling

### v1

@@ 6.4 Introduction to Do-Calculus
Judea Pearl’s do-calculus is a mathematical framework designed to infer causal relationships and predict outcomes based on interventions in complex systems. Central to Pearl’s causal inference is the concept of causal diagrams, or causal graphs, which represent variables as nodes and causal relationships as directed edges between them. These causal diagrams enable formal analysis of causality by allowing us to reason about the effects of interventions [pearl_book, Pearl1995]. The do-calculus itself consists of rules that help transform expressions involving interventions into expressions based solely on observable quantities, thus making causal inferences feasible even from observational data [Pearl2010].
@@ 6.4.1 Back-Door Criterion and Causal Inference
This criterion is critical in adjusting for confounding variables and obtaining unbiased estimates of causal effects [Pearl1995, Pearl2010].
@@ 6.4.2 Rules of Do-Calculus
These rules provide a systematic approach for transforming expressions involving causal interventions, allowing researchers to determine whether a causal effect can be estimated from observational data. The rules also facilitate the derivation of expressions that capture the direct effect of an action, independent of confounding variables [pearlcauses].

### v2

@@ 6.4 Introduction to Do-Calculus
Judea Pearl’s do-calculus is a mathematical framework designed to infer causal relationships and predict outcomes based on interventions in complex systems. Central to Pearl’s causal inference is the concept of causal diagrams, or causal graphs, which represent variables as nodes and causal relationships as directed edges between them. These causal diagrams enable formal analysis of causality by allowing us to reason about the effects of interventions [32, 31]. The do-calculus itself consists of rules that help transform expressions involving interventions into expressions based solely on observable quantities, thus making causal inferences feasible even from observational data [33].
@@ 6.4.1 Back-Door Criterion and Causal Inference
This criterion is critical in adjusting for confounding variables and obtaining unbiased estimates of causal effects [31, 33].
@@ 6.4.2 Rules of Do-Calculus
These rules provide a systematic approach for transforming expressions involving causal interventions, allowing researchers to determine whether a causal effect can be estimated from observational data. The rules also facilitate the derivation of expressions that capture the direct effect of an action, independent of confounding variables [11].